OpenStudy (anonymous):

A coin is tossed three times. (a) What is the probability that the coin will land tails at least twice?

OpenStudy (anonymous):

n trial = 3 p (tail ) = 1/2, q = 1/2 at least twice = P( 2 tails ) + P ( 3 tails) Can you apply binomial theorem?